In a certain country, there is a bank failure once every six days, on average. What is the probability of at least one bank failure in the next 30 days?
Is this set up 1 less 1/6?

Answer by khwang(438)

I think youu have to use Poisson distribution
the mean (avg) of one bank failing is 1 for every 6 days.
Hence, the mean is 5 =30/6 banks failing for 30 days .
Now lambda = 5,
Let the random variable of Poisson dist. with mean u = 5 (in 30 days).

We have Pr(X=k) = u^k/k! *e^(-u)
Hence, Pr(X =>1) = 1- Pr(X =0) = 1 - e^(-5) = 0.993
Kenny